# Valentina Bais and Rafael Torres Scuola Internazionale Superiori di Studi Avanzati (SISSA) Via Bonomea 265 34136

### v2

@@ 3.3. Proof of Theorem C
The four-manifold $\mathcal{R}_{2p}$ was constructed in (3.1) for any odd $p\in\mathbb{Z}_{>0}$ and its homeomorphism type was pinned down in Lemma 11; see Remark 12. The claim that there does not exist a diffeomorphism $\mathcal{R}_{2p}\rightarrow\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ follows from Proposition 14. Indeed, if any such diffeomorphism between $\mathcal{R}_{2p}$ and $\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}}$ were to exist, it would induce a diffeomorphism between the degree-two coverings
$\mathcal{E}_{p}\rightarrow(({\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}})\setminus\nu(\gamma))\cup S_{p}=({\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathbb{P}}^{2}})\mathbin{\#}\Sigma_{p}$
@@
4.1. Seiberg-Witten invariants of smoothly reducible double coverings
The following proposition is a major trait in our proof of Theorem D. We direct the reader towards [34, Section 2.4], [8, Section 2.5] and the references there for the necessary background on Seiberg-Witten theory.
@@ 4.1. Seiberg-Witten invariants of smoothly reducible double coverings
(Kotschick-Morgan-Taubes [27, Proposition 2], Baykur-Stipsicz-Szabó [8, Corollary 2.11]). Let $X_{1}$ and $X_{2}$ be closed oriented smooth four-manifolds and suppose that $b_{1}(X_{2})=0=b_{2}^{+}(X_{2})$. There is a $\operatorname{Spin}^{\mathbb{C}}$-structure $(X,\mathfrak{s}_{X})$ on the connected sum $X=X_{1}\mathbin{\#}X_{2}$ that extends a $\operatorname{Spin}^{\mathbb{C}}$-structure $(X_{1},\mathfrak{s}_{X_{1}})$ such that
@@
Baykur-Stipsicz-Szabó observed in [8, Corollary 2.11] that Proposition 18 is a special case of the Seiberg-Witten blow-up formula [13]. Proposition 19, along with the blow up formula [13] (see [8, Proposition 2.8]). Proposition 18 will be used to argue the conclusion of the irreduciblity property in the statement of Theorem D, together with an argument due to Baykur-Stipsicz-Szabó [8].
@@
(cf. [8, Proof of Proposition 6.2]). Let $\widetilde{M}$ be a closed smooth oriented simply connected four-manifold and assume it satisfies at least one of the following two properties.
$\bullet$ The four-manifold $\widetilde{M}$ has exactly two Seiberg-Witten basic classes ±$\pm K_{\widetilde{M}}\in\mathfrak{B}_{\widetilde{M}}$ or
$\bullet$ The four-manifold $\widetilde{M}$ is minimal and it has several Seiberg-Witten basic classes {±$\{\pm K^{1}_{\widetilde{M}},\ldots,\pm K^{i}_{\widetilde{M}}\}\in\mathfrak{B}_{\widetilde{M}}$ with $i\geq 2$.
@@
is irreducible, provided $p$ is an odd integer. Furthermore, there is a $\operatorname{Spin}^{\mathbb{C}}$-structure $(\widehat{M}(p),\mathfrak{s}_{\widehat{M}(p)})$ on the degree-two covering
@@
corresponding to the subgroup $\mathbb{Z}/p\subset\pi_{1}(M(2p))=\mathbb{Z}/2p$ for which the value of the Seiberg-Witten invariant satisfies
(4.1) $\operatorname{SW}_{\widehat{M}(p)}(\mathfrak{s}_{\widehat{M}(p)})=\operatorname{SW}_{\widetilde{M}\mathbin{\#}\Sigma_{p}}(\mathfrak{s_{\mathbin{\#}}})=\operatorname{SW}_{\widetilde{M}}(\mathfrak{s}_{\widetilde{M}})$
for $\operatorname{Spin}^{\mathbb{C}}$-structures $(\widetilde{M}\mathbin{\#}\Sigma_{p},\mathfrak{s}_{\mathbin{\#}})$ and $(\widetilde{M},\mathfrak{s}_{\widetilde{M}})$.
The reader might have already noticed that the properties that the simply connected four-manifold $\widetilde{M}$ of Proposition 19 is required to have imply that it is irreducible.
@@
where neither of the four-manifolds $M_{1}$ and $M_{2}$ is a homotopy four-sphere. Since the finite cyclic group $\pi_{1}(M(2p))=\mathbb{Z}/2p$ does not decompose as a free product of two non-trivial groups, without loss of generality we can assume that
$\pi_{1}(M_{1})=\mathbb{Z}/2p\quad\text{and}\quad\pi_{1}(M_{2})=\{1\}$
and $b_{2}(M_{2})\geq 1$. We hence obtain a decomposition of the degree-two covering
(4.3) $\widehat{M}(p)=\widehat{M}_{1}\mathbin{\#}2M_{2}$
from (4.2) where $\widehat{M}_{1}$ is the degree-two covering of $M_{1}$. Since by Lemma 8 there is an identification $\widehat{M}(p)=\widetilde{M}\mathbin{\#}\Sigma_{p}$, Proposition 18 implies that there is a $\operatorname{Spin}^{\mathbb{C}}$-structure $(\widehat{M}(p),\mathfrak{s}_{\widehat{M}(p)})$ with $\operatorname{SW}_{\widehat{M}(p)}(\mathfrak{s}_{\widehat{M}(p)})\neq 0$ and it follows that $b_{2}^{+}(M_{2})=0$. In particular, the intersection form of $M_{2}$ is isomorphic to $b_{2}(M_{2})\langle-1\rangle.$
At this point, we employ an argument due to Baykur-Stipsicz-Szabo [8] to prove Proposition 19. Consider the map
@@
assigning to a given $\operatorname{Spin}^{\mathbb{C}}$-structure $(\widehat{M}(p),\mathfrak{s})$ its first Chern class with $\mathbb{Q}$-coefficients. Suppose first that the four-manifold $\widetilde{M}$ has exactly two Seiberg-Witten basic classes. As argued by Baykur-Stipsicz-Szabó [8, Proof of Proposition 6.2], the blow up formula [13] (see [8, Proposition 2.8]) and Proposition 18 imply that the image of the set
(4.5) $\{\mathfrak{s}_{\widehat{M}(p)}\in\operatorname{Spin}^{\mathbb{C}}(\widehat{M}(p)):\operatorname{SW}_{\widehat{M}(p)}(\mathfrak{s}_{\widehat{M}(p)})\neq 0\}$
under the map (4.4) has two elements, given that the four-manifold $\widetilde{M}$ is assumed to have exactly two Seiberg-Witten basic classes and the covering decomposes as $\widehat{M}(p)=\widetilde{M}\mathbin{\#}\Sigma_{p}$ by Lemma 8. On the other hand, the decomposition (4.3) implies that the image of the set (4.5) under (4.4) contains at least four elements. This is a contradiction and we conclude that the four-manifold $M(2p)$ is irreducible. A slight extension of Baykur-Stipsicz-Szabó’s argument allows us to reach a similar contradiction and conclude that $M(2p)$ is irreducible under the assumptions that the four-manifold $\widetilde{M}$ admits several Seiberg-Witten basic classes and that it is minimal. In this case, the image of the set (4.5) under the map (4.4) has $i$ elements, where $i>2$ is the number of Seiberg-Witten basic classes of $\widetilde{M}$. If there were indeed a decomposition as in (4.3), then the number of elements of the image of the set (4.5) under the map (4.4) would be strictly larger than $i$. We conclude that the four-manifold $M(2p)$ is irreducible. The claim on the value (4.1) of the Seiberg-Witten invariant is an immediate consequence of Proposition 18.
@@ 4.2. Proof of Theorem D
Denote by $X_{n}(2)\mathrel{\mathop{:}}=\widetilde{X}_{n}/\tau_{n}$ the closed smooth oriented four-manifold with fundamental group of order two that arises as the quotient of the closed oriented simply connected four-manifold $\widetilde{X}_{n}$ by its orientation-preserving fixed-point free involution $\tau_{n}:\widetilde{X}_{n}\rightarrow\widetilde{X}_{n}$. The classification result of Hambleton-Kreck [21, Theorem C] says that $X_{n}(2)$ is homeomorphic to $Y\mathbin{\#}\Sigma_{2,k}$ for a given closed simply connected topological four-manifold $Y$ with Kirby-Siebenmann invariant $\operatorname{KS}(Y)=k=\Sigma_{2,k}$. Fix an odd positive integer $p\in\mathbb{Z}_{>0}$ and assemble
@@
and $S_{p}$ is the compact four-manifold that was manufactured in Section 2.1. Lemma 6 says that the fundamental group of the four-manifold (4.6) is the finite cyclic group $\pi_{1}(X_{n}(2p))=\mathbb{Z}/2p$. Moreover, there is a homeomorphism between $X_{n}(2p)$ and $Y\mathbin{\#}\Sigma_{2p,k}$ by Corollary 5 (and [21, Theorem C] as well).
Lemma 8 says that the degree-two covering $X_{n}(p)$ that corresponds to the subgroup $\mathbb{Z}/p\subset\mathbb{Z}/2p$ is diffeomorphic to $X_{n}(p)=\widetilde{X}_{n}\mathbin{\#}\Sigma_{p}$. Proposition 19 says that the four-manifold (4.6) is irreducible. At this point, we have proven all clauses of Theorem D except for the third one. To conclude on the existence of infinitely many pairwise non-diffeomorphic irreducible smooth structures on $Y\mathbin{\#}\Sigma_{2p,k}$ and conclude the proof of the theorem, we employ Lemma 9 and distinguish them in terms of their double coverings: the value of the Seiberg-Witten function is assumed to be $\operatorname{SW}_{\widetilde{X}_{n}}(\pm K_{n})=\pm n^{2}$ on its sole pair of basic classes and $\widetilde{X}_{n_{1}}$ is diffeomorphic to $\widetilde{X}_{n_{2}}$ if and only if $n_{1}=n_{2}$. Notice that this argument is due to Stipsicz-Szabó [34] and Baykur-Stipsicz-Szabó [8, Proposition 2.14]. $\square$
@@
Theorem D and the following result are the main ingredients in our proof of Theorem E.
@@
There is a closed smooth irreducible simply connected four-manifold $\widetilde{X}_{m,n}$ with
@@
A substantial region of Theorem 20 is due to Arabadji-Morgan’s extensive study on the existence of fixed-point free orientation-preserving involutions on the closed irreducible symplectic simply connected four-manifolds constructed in [2, 19]. Theorem D follows from Theorem 20 and the last clause of Theorem D. $\square$
Remark 21.
As Porter Morgan kindly pointed out to us, several of the irreducible simply-connected four-manifolds with an involution constructed in [4] have more than two Seiberg-Witten basic classes. A stronger conclusion of infinitely many exotic smooth structures in Theorem E using our mechanism would depend on an improvement of Arabadji-Morgan’s result in this direction.

### v3

@@ 3.3. Proof of Theorem C
The four-manifold $\mathcal{R}_{2p}$ was constructed in (3.1) for any odd $p\in\mathbb{Z}_{>0}$ and its homeomorphism type was pinned down in Lemma 11. The claim that there does not exist a diffeomorphism $\mathcal{R}_{2p}\rightarrow\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}}$ follows from Proposition 13. Indeed, if any such diffeomorphism between $\mathcal{R}_{2p}$ and $\Sigma_{2p}\mathbin{\#}4\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}}$ were to exist, it would induce a diffeomorphism between the degree-two covers
$\mathcal{E}_{p}\rightarrow(({\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}})\setminus\nu(\gamma))\cup S_{p}=({\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}\mathbin{\#}9\overline{{\mathbb{C}\mkern-0.5mu\mathrm{P}}^{2}})\mathbin{\#}\Sigma_{p}$
@@
4.1. Seiberg-Witten invariants of smoothly reducible double covers
Let $\operatorname{Spin}^{\mathbb{C}}(X)$ be the set of $\operatorname{Spin}^{\mathbb{C}}$-structures on a closed oriented 4-manifold $X$. We regard the Seiberg-Witten invariant of $X$ as a function
$\operatorname{SW}_{X}:\operatorname{Spin}^{\mathbb{C}}(X)\rightarrow\mathbb{Z}.$
A cohomology class $K\in H^{2}(X;\mathbb{Z})$ is a Seiberg-Witten basic class if $\operatorname{SW}_{X}(\mathfrak{s})\neq 0$ for a $\operatorname{Spin}^{\mathbb{C}}$-structure $(X,\mathfrak{s})$ with $c_{1}((X,\mathfrak{s})=K$. The set of Seiberg-Witten basic classes of the four-manifold $X$ is denoted by $\mathfrak{B}_{X}$. We direct the reader towards [8, 21, 40, 42] and the references there for the necessary background on Seiberg-Witten theory.
The proofs of our main results require us to determine the Seiberg-Witten invariants of connected sums of four-manifolds $X=X_{1}\mathbin{\#}X_{2}$. These invariants are known to be determined in terms of the summands $X_{1}$ and $X_{2}$ of the connected sum $X$. We start this section by recalling a pair of results for such purpose. For any $\operatorname{Spin}^{\mathbb{C}}$-structure $(X_{1},\mathfrak{s}_{X_{1}})$, denote its extension to $X=X_{1}\mathbin{\#}X_{2}$ by $(X,\mathfrak{s}_{X})$. Fintushel-Stern’s blow up formula [14] establishes a relation between the invariants of $X$ and $X_{1}$.
@@ 4.1. Seiberg-Witten invariants of smoothly reducible double covers
[14]; cf. [8, Proposition 2.8]. Let $\{(X_{i},\mathfrak{s}_{i}):i=1,2\}$ be $\operatorname{Spin}^{\mathbb{C}}$-structures where the underlying four-manifolds are closed and oriented. Consider nonnegative integers $\{n_{i}:i=1,\ldots,k\}$ given by $\langle c_{1}(\mathfrak{s}_{2}),e_{i}\rangle=2n_{i}+1$ that satisfy the identity
$\frac{1}{4}(c_{1}(\mathfrak{s}_{1}^{2}-2\chi(X_{1})-3\sigma(X_{1}))-\overset{k}{\underset{i=1}{\sum}}n_{i}(n_{i}+1)\geq 0.$
The Seiberg-Witten invariants satisfy
$\operatorname{SW}_{X_{1}}(\mathfrak{s}_{1})=\pm\operatorname{SW}_{X_{1}\mathbin{\#}X_{2}}(\mathfrak{s}_{1}\mathbin{\#}\mathfrak{s}_{2}).$
Baykur-Stipsicz-Szabó observed that the next proposition is a special case of the Seiberg-Witten blow up formula in [8, Corollary 2.11]. The result was also obtained by Kotschick-Morgan-Taubes [30, Proposition 2].
Proposition 21.
([8, 30]). Let $X_{1}$ and $X_{2}$ be closed oriented four-manifolds and suppose that $b_{1}(X_{2})=0=b_{2}^{+}(X_{2})$. For any $\operatorname{Spin}^{\mathbb{C}}$-structure $(X_{1},\mathfrak{s}_{X_{1}})$, denote by $(X,\mathfrak{s}_{X})$its extension to the connected sum $X=X_{1}\mathbin{\#}X_{2}$. The corresponding Seiberg-Witten invariants satisfy
@@
The next proposition is the main technical contribution in this section. Besides the pair of proposition stated before, we employ arguments of Baykur-Stipsicz-Szabó [8, Proof of Proposition 6.2] and of Park [37, Proof of Theorem 6] in its proof.
@@
Let $\widetilde{M}$ be a closed oriented simply connected four-manifold that satisfies one of the following two properties.
$\bullet$ It has exactly two Seiberg-Witten basic classes $\pm K_{\widetilde{M}}\in\mathfrak{B}_{\widetilde{M}}$.
[…]
@@
is irreducible provided $p\in\mathbb{Z}_{>0}$ is odd. Furthermore, there is a $\operatorname{Spin}^{\mathbb{C}}$-structure $(\widehat{M}(p),\mathfrak{s}_{\widehat{M}(p)})$ on the degree-two cover
@@
that corresponds to the subgroup $\mathbb{Z}/p\subset\pi_{1}(M(2p))=\mathbb{Z}/2p$ for which the value of the Seiberg-Witten invariant satisfies
[…]
A symplectic four-manifold is minimal if and only if it contains no embedded two-sphere of self-intersection minus one [21, Remark 10.2.5 (a)]. The reader might have already noticed that the properties that the simply connected four-manifold $\widetilde{M}$ of Proposition 22 is required to have imply that it is irreducible [42, 29, 22].
@@
where neither of the four-manifolds $M_{1}$ and $M_{2}$ is a homotopy four-sphere. Since the finite cyclic group $\pi_{1}(M(2p))=\mathbb{Z}/2p$ does not decompose as a free product of two non-trivial groups, without loss of generality we can assume that $\pi_{1}(M_{1})=\mathbb{Z}/2p$, $\pi_{1}(M_{2})=\{1\}$ and $b_{2}(M_{2})\geq 1$. This yields a decomposition of the degree-two cover of (4.2) as
(4.3) $\widehat{M}(p)=\widehat{M}_{1}\mathbin{\#}2M_{2},$
where $\widehat{M}_{1}$ is the degree-two cover of $M_{1}$. We also have the decomposition
(4.4) $\widehat{M}(p)=\widetilde{M}\mathbin{\#}\Sigma_{p}$
from Lemma 8.
These identifications and Proposition 21 imply that there are bijections
(4.5) $\mathfrak{B}_{\widehat{M}_{1}\mathbin{\#}2M_{2}}\rightarrow\mathfrak{B}_{\widetilde{M}\mathbin{\#}\Sigma_{p}}\rightarrow\mathfrak{B}_{\widetilde{M}}$
among the Seiberg-Witten basic classes. If $\widetilde{M}$ has non-trivial Seiberg-Witten invariant, then $\widehat{M}(p)$ also has non-trivial Seiberg-Witten invariant and $b_{2}^{+}(2M_{2})=0$ [21, Theorem 2.4.6] (this is the case in both instances of the proposition ([42])). This implies that the intersection form of $Q_{2M_{2}}$ is negative-definite and Donaldson’s theorem [12] says that $Q_{2M_{2}}\cong 2b_{2}(M_{2})\langle-1\rangle$. Let $\{e^{1}_{1},e^{2}_{1},\ldots,e^{1}_{b_{2}(M_{2})},e^{2}_{b_{2}(M_{2})}\}$ be a basis for the second cohomology group $H^{2}(2M_{2};\mathbb{Z})=\mathbb{Z}^{2b_{2}(M_{2})}$ with $(e^{i}_{j})^{2}=-1$ for each $i=1,2$ and $j=1,\ldots,b_{2}(M_{2})$.
From (4.3) and using the neck pinching argument [12, 29], we conclude that there is at least one basic class $K_{\widehat{M}_{1}}\in\mathcal{B}_{\widehat{M}_{1}}$ and the Seiberg-Witten invariant is $\operatorname{SW}_{\widehat{M}_{2}}(\mathfrak{s}_{\widehat{M}_{1}})\neq 0$ for $c_{1}(\mathfrak{s}_{\widehat{M}_{1}})=K_{\widehat{M}_{1}}$. The set of basic classes $\mathfrak{B}_{\widehat{M}_{1}\mathbin{\#}2M_{2}}$ consists of elements of the form
(4.6) $K_{\widehat{M}_{1}}\pm e^{1}_{1}\pm e^{2}_{2}\pm\cdots\pm e^{1}_{b_{2}(M_{2})}\pm e^{2}_{b_{2}(M_{2})}$
for $K_{\widehat{M}_{1}}\in\mathfrak{B}_{\widehat{M}_{1}}$ [21, Theorem 2.4.10, Remark 2.4.11]. At this point, we employ an argument due to Baykur-Stipsicz-Szabó [8] to prove the first instance of Proposition 22. Consider the map
@@
assigning to a given $\operatorname{Spin}^{\mathbb{C}}$-structure $(\widehat{M}(p),\mathfrak{s})$ its first Chern class with $\mathbb{Q}$-coefficients. The punchline is that the number of elements of the image of the set
(4.8) $\{\mathfrak{s}\in\operatorname{Spin}^{\mathbb{C}}(\widehat{M}(p)):\operatorname{SW}_{\widehat{M}(p)}(\mathfrak{s})\neq 0\}$
under the map (4.7) differs for the decompositions (4.3) and (4.4), and this yields a contradiction. Suppose the four-manifold $\widetilde{M}$ has exactly two Seiberg-Witten basic classes. Proposition 20 and Proposition 21 along with the decomposition (4.4) imply that the image of the set (4.8) under the map (4.7) has two elements, given that the four-manifold $\widetilde{M}$ is assumed to have exactly two Seiberg-Witten basic classes. On the other hand, the decomposition (4.3) implies that the image of the set (4.8) under (4.7) contains more than two elements by (4.6). This is a contradiction and we conclude that the four-manifold $M(2p)$ is irreducible.
We now address the second instance of the proposition using an argument due to Park [37, Proof of Theorem 6] to finish the proof. Suppose that $\widetilde{M}$ admits a symplectic structure and that it is minimal. The previous arguments show that there is an isomorphism
(4.9) $\xi:H^{2}(\widetilde{M};\mathbb{Z})\rightarrow H^{2}(\widehat{M}_{1}\mathbin{\#}2M_{2};\mathbb{Z})$
that preserves the cup product pairing and that restricts to a bijection $\mathfrak{B}_{\widetilde{M}}\rightarrow\mathfrak{B}_{\widehat{M}_{1}\mathbin{\#}2M_{2}}$. Taubes showed that $c_{1}(\widetilde{M},\omega_{\widetilde{M}})\in\mathfrak{B}_{\widetilde{M}}$ and its image under (4.9) is $\xi(c_{1}(\widetilde{M},\omega_{\widetilde{M}}))\in\mathfrak{B}_{\widehat{M}_{1}\mathbin{\#}2M_{2}}$, where the elements of the latter set are of the form (4.6). Without loss of generality, we can assume that $\xi(c_{1}(\widetilde{M},\omega_{\widetilde{M}}))=K_{\widehat{M}_{1}}-e^{1}_{1}-e_{1}^{2}-\cdots-e^{2}_{b_{2}(M_{2})}-e^{2}_{b_{2}(M_{2})}$. In particular,
(4.10) $\xi(c_{1}(\widetilde{M},\omega_{\widetilde{M}}))+2e_{1}^{1}=K_{\widehat{M}_{1}}+e_{1}^{1}-e_{1}^{2}-\cdots-e^{2}_{b_{2}(M_{2})}\in\mathfrak{B}_{\widehat{M}_{1}\mathbin{\#}2M_{2}}.$
[…]
Remark 23.
The previous modification of the argument Park yields a proof of the following result, which is a slight generalization of [37, Theorem 6].
Theorem 24.
Let $Z$ be a closed minimal symplectic four-manifold with $b_{2}^{+}(Z)>1$ and let $M$ be a closed four-manifold with cyclic fundamental group $\pi_{1}(M)=\mathbb{Z}/q$. Suppose that there is an isomorphism
(4.11) $\xi:H^{2}(Z)\rightarrow H^{2}(M)$
that preserves the cup product and that restricts to a one-to-one correspondence $\mathfrak{B}_{Z}\rightarrow\mathfrak{B}_{M}$. The four-manifold $M$ is irreducible.
@@ 4.2. Proof of Theorem D
Denote by $X_{n}(2)\mathrel{\mathop{:}}=\widetilde{X}_{n}/\tau_{n}$ the closed smooth oriented four-manifold with fundamental group of order two that arises as the quotient of the closed oriented simply connected four-manifold $\widetilde{X}_{n}$ by its orientation-preserving fixed-point free involution $\tau_{n}:\widetilde{X}_{n}\rightarrow\widetilde{X}_{n}$. The classification result of Hambleton-Kreck [23, Theorem C] says that $X_{n}(2)$ is homeomorphic to $Y\mathbin{\#}\Sigma_{2,k}$ for a given closed simply connected topological four-manifold $Y$ with Kirby-Siebenmann invariant $\operatorname{KS}(Y)=k=\Sigma_{2,k}$. Fix an odd positive integer $p\in\mathbb{Z}_{>0}$ and assemble
@@
and $S_{p}$ is the compact four-manifold that was manufactured in Section 2.1. Lemma 6 says that the fundamental group of the four-manifold (4.12) is the finite cyclic group $\pi_{1}(X_{n}(2p))=\mathbb{Z}/2p$. Moreover, there is a homeomorphism between $X_{n}(2p)$ and $Y\mathbin{\#}\Sigma_{2p,k}$ by Corollary 5 (and [23, Theorem C] as well).
Lemma 8 says that the degree-two cover $X_{n}(p)$ that corresponds to the subgroup $\mathbb{Z}/p\subset\mathbb{Z}/2p$ is diffeomorphic to $X_{n}(p)=\widetilde{X}_{n}\mathbin{\#}\Sigma_{p}$. Proposition 22 says that the four-manifold (4.12) is irreducible. At this point, we have proven all clauses of Theorem D except for the third one. To conclude on the existence of infinitely many pairwise non-diffeomorphic irreducible smooth structures on $Y\mathbin{\#}\Sigma_{2p,k}$ and conclude the proof of the theorem, we employ Lemma 9 and distinguish them in terms of their double covers as done in [40]: the value of the Seiberg-Witten function is assumed to be $\operatorname{SW}_{\widetilde{X}_{n}}(\pm\mathfrak{s}_{n})=\pm n^{2}$ and there is only one pair of basic classes. Hence, $\widetilde{X}_{n_{1}}$ is diffeomorphic to $\widetilde{X}_{n_{2}}$ if and only if $n_{1}=n_{2}$. $\square$
@@
Theorem D and the following result are the main ingredients in our proof of Theorem E. The case $m=1$ is completely covered in [8, Theorem 1.1].
@@
There is a closed irreducible simply connected symplectic four-manifold $\widetilde{X}_{m,n}$ with
@@
Theorem E follows from Theorem 25 and Theorem D. $\square$